OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot.

OpenStudy (campbell_st):

206 = pir^2 206/pi = r^2 \[r = \sqrt{206/pi}\]